OpenStudy (anonymous):

lim(n->inf) [sqrt(n)*(sqrt(3n+1) - sqrt(3n))] any help is appreciated! Thanks.

OpenStudy (dumbcow):

distribute \[\sqrt{3n^{2}+n}-\sqrt{3n^{2}}\] \[= \sqrt{3n^{2}+n} -\sqrt{3} n\] sub n = 1/u , change limit to u->0 \[\lim_{u \rightarrow 0} \sqrt{\frac{3}{u^{2}}+\frac{1}{u}}- \frac{\sqrt{3}}{u}\] simplify \[\lim_{u \rightarrow 0} \frac{\sqrt{3+u} -\sqrt{3}}{u}\]

OpenStudy (dumbcow):

this results in 0/0 you can now either use L'hopitals rule or multiply top/bottom by conjugate

OpenStudy (zarkon):

use the definition of the derivative

OpenStudy (zarkon):

that is what you have

OpenStudy (zarkon):

\[f(x)=\sqrt{x}\] the limit is \[f'(3)\]

OpenStudy (anonymous):

@dumbcow, @zarkon can i ask how you knew to look for that relationship? I see it now, obviously, after you pointed it out.. but I would have never seen that otherwise.

OpenStudy (anonymous):

also, i wish i could give both of you the medal, but since dumbcow was here first, it's only fair.. sorry zarkon!

OpenStudy (zarkon):

I'll live ;)